A rectangular fish tank is 2 feet high and has a volume of 48 feet cubed. What could be a possible set of whole number dimensions for the base of the tank?
Answer by Theo(13342)

volume = height * area of base.

let v = volume, h = height, a = area of base.

v = h*a

solve for a to get:

a = v/h

v = 48
h = 2

a = 48/2 = 24

the area of the base is 24 square inches.

24 = 1*24 = 2*12 = 3*8 = 4*6

possible dimensions for the base where the dimensions are in whole numbers are:


                   length          width
                     24              1
                     12              2
                      8              3
                      6              4